OpenStudy (anonymous):

prove the identity: 1 + cos 2(x)/ sin 2(x) = cot(x)

OpenStudy (anonymous):

Please help me! :)

OpenStudy (dumbcow):

first use double angle identities \[\sin 2x = 2 \sin x \cos x\] \[\cos 2x = \cos^2 x - \sin^2 x\] \[\rightarrow \frac{1+ \cos^2 x -\sin^2 x}{2 \sin x \cos x}\] From pythagorean identity \[\cos^2 x = 1-\sin^2 x\] \[\rightarrow \frac{2 \cos^2 x}{2 \sin x \cos x}\] i think you can simplify it from here

OpenStudy (anonymous):

I really appreciate this, however do you mind finishing the problem out? I don't understand how you will get cot(x) from this?

OpenStudy (anonymous):

Factor out 2cos(x).

OpenStudy (dumbcow):

what cancels?

OpenStudy (anonymous):

so are you left with cos/2sinx ?

OpenStudy (anonymous):

wait i meant cos/sinx

OpenStudy (anonymous):

ooohh I think I get it and cot x = cos x/ sin x, right!!!???

OpenStudy (anonymous):

Yes!

OpenStudy (dumbcow):

yes correct :) always keep a trig reference chart around to help with identities and what cot is

OpenStudy (anonymous):

Whoo hoo! :)

OpenStudy (anonymous):

ok - thanks guys!

OpenStudy (anonymous):

I do have one more question.... can you help?

OpenStudy (anonymous):

if log(subscript a) x=7, what is the value of log(subscript a) (a/x^2) ?

OpenStudy (dumbcow):

-13